OpenStudy (anonymous):

5. What is the value of tan 45°? (Points : 1) Option A: √2/2 Option B: √3/2 Option C: 1 Option D: √3

OpenStudy (anonymous):

this one is diffrent b/c the last one asked for the cos and this one asks for the tan

OpenStudy (anonymous):

any clues?

OpenStudy (anonymous):

no - sory im utterly an official failer when it comes to this >.>

OpenStudy (anonymous):

find 45 degrees on the unit circle of the last page of the attached cheat sheet

OpenStudy (anonymous):

you will see a coordinate, it should look like this \[(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2})\]

OpenStudy (anonymous):

the first coordinate is cosine, so \(\cos(45)=\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

the second coordinate is sine, so \[\sin(45)=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

right but im not sure what the tan is thats were i get confused

OpenStudy (anonymous):

and finally the tangent is \(\frac{\sin(x)}{\cos(x)}\) in this case since they are the same, you get \[\tan(45)=1\]

OpenStudy (anonymous):

oh i see cool ^.^ this makes sence but im stuck on a diffrent question could you help me, im not sure if it goes along with this way of mathmatical term.

OpenStudy (anonymous):

sure go ahead and ask

OpenStudy (anonymous):

On a baseball diamond,1st base, 2nd base, 3rd base, and home plate form a square. If the ball is thrown from 1st base to 2nd base and then from 2nd base to home plate, how many feet has the ball been thrown? (Points : 1) Option A: 90√2 Option B: 180√2 Option C: 90+90√2 Option D: 270√2

OpenStudy (anonymous):

the diagonal from second to home is merely the hypotenuse of a right triangle with equal sides of 90 ft. use the Pythagorean Theorem for a square: s^2 + s^2 = d^2 solving for d: 2s^2=d^2 or d=[2s^2]^1/2 or d=1.414s okay, the ball goes 90 feet from 1st to 2nd, and 1.414(90 feet) from 2nd to home. that's a total of 217.3 feet.

OpenStudy (anonymous):

first to second is \(90\) then second to home is the diagonal of a square with side \(90\) so that length is \(90\sqrt{2}\) and your answer is \[90+90\sqrt{2}\]

OpenStudy (anonymous):

thank you so so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

keep that cheat sheet, it is very useful for this trig stuff